OpenStudy (anonymous):

First factor the trinomial as a binomial square and then use the square root method to solve. Round the solution to the nearest tenth. x2 - 14x + 49 = 3

OpenStudy (anonymous):

Just follow the instructions.\[x^2 - 14x + 49 = 3\implies (x-7)^2=3 \implies x-7= \pm \sqrt3 \implies x=7 \pm \sqrt3\]Note that the LHS of the equation is a perfect square.